OpenStudy (toxicwaste):

What is the slope of the graph of 6x + 12y = 7? A. -2 B. -1/2 C. 1/2 D. 2

OpenStudy (mathstudent55):

Solve for y.

jagr2713 (jagr2713):

Use the formula Y=Mx+b

OpenStudy (mathstudent55):

\(Ax + By = C\) \(By = -Ax + C\) \(y = -\dfrac{A}{B}x + \dfrac{C}{B} \) The slope of a line whose equation is in standard form is \(-\dfrac{A}{B} \)

OpenStudy (toxicwaste):

so the answer would be -1/2

jagr2713 (jagr2713):

Correct :D

OpenStudy (toxicwaste):

:D

OpenStudy (mathstudent55):

Yes. 6x + 12y = 7 12y = - 6x + 7 y = -(6/12)x + 7/12 y = -(1/2)x + 7/12 Slope = -1/2

OpenStudy (toxicwaste):

thx u @jagr2713 and @mathstudent55

jagr2713 (jagr2713):

Welcome